# Sim2real Image Translation Enables Viewpoint-Robust Policies from Fixed-Camera Datasets

### v1

@@ Abstract
Vision-based policies for robot manipulation have achieved significant recent success, but are still brittle to distribution shifts such as camera viewpoint variations. Robot demonstration data is scarce and often lacks appropriate variation in camera viewpoints. Simulation offers a way to collect robot demonstrations at scale with comprehensive coverage of different viewpoints, but presents a visual sim2real challenge. To bridge this gap, we propose MANGO – an unpaired image translation method with a novel segmentation-conditioned InfoNCE loss, a highly-regularized discriminator design, and a modified PatchNCE loss. We find that these elements are crucial for maintaining viewpoint consistency during sim2real translation. When training MANGO, we only require a small amount of fixed-camera data from the real world, but show that our method can generate diverse unseen viewpoints by translating simulated observations. In this domain, MANGO outperforms all other image translation methods we tested. Imitation-learning policies trained on data augmented by MANGO are able to achieve success rates as high as 60% on views that the non-augmented policy fails completely on. For more results, visit: https://sites.google.com/view/sim2real-viewpoints.
@@ I Introduction
Figure 1: (a) Standard image translation methods fail to generalize to new viewpoints when trained on fixed-viewpoint target domain dataset. (b) Our method, MANGO, enables realistic generation of unseen viewpoints, which are used to improve the robustness of downstream robot polices.
For example, many tabletop manipulation datasets employ fixed, third person cameras for observations [12, 39, 23, 41, 21, 6, 46], making downstream policies brittle to camera viewpoint shifts. Indeed, we have observed that when robot polices are trained on fixed-camera datasets, changes in camera viewpoint during deployment cause success rates to crater to zero (Table IV). Cameras are often fixed during demo collection to ensure consistent visual observations, avoid repeated calibration with depth or motion capture sensors, or simply for convenience. Generalizing to truly unseen viewpoints is difficult because a change in viewpoint affects the entire scene, and the model must implicitly estimate the robot’s position relative to the new camera position.
@@ III-A Image-to-image Translation
We propose MANGO, a novel unpaired image-to-image translation method to translate visual observations from sim to real (Figure 2). MANGO is trained on on a small, fixed-viewpoint dataset of real images yet is capable of accurately translating viewpoint-diverse observations from a simple digital twin to realistic unseen viewpoints.
The objective of unpaired image-to-image translation is to translate images from domain $A$ to domain $B$ without access to a paired dataset of images $\mathcal{D}_{paired}=\{d_{A},\;d_{B}|d_{A}\in A,\;d_{b}\in B\}_{i=0}^{N}$. Instead, we learn from two disjoint datasets $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$, where our domain $A$ is simulation, our domain $B$ is the real-world, and $|\mathcal{D}_{A}|>|\mathcal{D}_{B}|$. This problem is considered unsupervised because there is no label, or ground-truth image, in $\mathcal{D}_{B}$ that images in $\mathcal{D}_{A}$ should map to.
@@ III-B Style Loss
One assumption underlying image-translation GANs, such as CycleGAN[51] and CUT [29], is that the shared attributes among all images in $D_{B}$ constitute the target domain ”style”. However, our real-world robot image observations in $\mathcal{D}_{B}$ only differ from one another in robot and object poses. Much of the image content, such as the background and tabletop, is nearly identical in all images in $\mathcal{D}_{B}$. A naïve discriminator will memorize the repetitive details and force the generator to recreate them. To mitigate this problem, Pix2pix [17] proposed a ”PatchGAN”, where the discriminator only receives local image patches and cannot therefore enforce global image elements. We take this a step further and randomly sample patch locations and apply per-patch random rotations. This process is shown in Figure 2. The result is a highly-regularized discriminator capable of enforcing the style of $\mathcal{D}_{B}$ on images with viewpoints not seen in $\mathcal{D}_{B}$.
@@ III-C1 Modified PatchNCE Loss
The PatchNCE loss consists of an InfoNCE loss across encoder features generated by a source domain image $d_{A}\in D_{A}$ and its corresponding translated output image $G(d_{A})$. For an input image $d$, we randomly sample $N$ latent features from the encoder’s feature map at $L$ different layers. We call the set of features at layer $l$ $\mathcal{Z}_{l}$ and $|\mathcal{Z}_{l}|=N\ \forall\ l\in\{l_{0},...,l_{L}\}$. The translated image $\hat{d}$ is passed through the encoder again to obtain $|\hat{\mathcal{Z}}_{l}|\ \forall\ l\in\{l_{0},...,l_{L}\}$. All $\hat{\mathcal{Z}}_{l}$ are obtained from the same feature map indices as in $\mathcal{Z}_{l}$.
The InfoNCE loss for feature $i$ in encoder layer $l$ is given by Equation 2. This is the categorical cross-entropy loss on the probability that a feature $\mathbf{z}\in\mathcal{Z}$ will be correctly classified as the corresponding feature in $\hat{\mathcal{Z}}$, based on a scoring function $\rho_{l}(\cdot)$. See [29] for further details. $\tau$ is a temperature hyperparameter.
@@
The assumption behind Equation 2 is that input and output features from the same feature map locations are “positive” samples and should have high similarity scores. All other features are “negative” samples and should be repelled. However, we observe that many different input image patches are highly similar due to repeated patterns or textures in robot image datasets, which include background elements, the tabletop, etc. Furthermore, in the simulated image dataset $\mathcal{D}_{A}$, these regions all have identical pixel values. Therefore, Equation 2 will repel many false negative features.
@@ III-C2 Segmentation NCE Loss
We leverage the simulator used to generate $\mathcal{D}_{A}$ to obtain ground-truth segmentation maps for each image. We propose an InfoNCE loss which clusters generator features by segmentation category in order to ensure that object boundaries are preserved during translation. We note that for object-centric manipulation, this is a crucial aspect to preserve when training policies.
Each image in $\mathcal{D}_{A}$ contains $C$ segmentation classes and each feature $\mathbf{z}_{i}\in\mathcal{Z}$ generated from $d\sim D_{A}$ has an associated class label $y_{i}\in\mathcal{Y}$. In the case when a layer’s feature map is of a lower resolution than the input image, we scale the image segmentation with nearest-neighbors downsampling to obtain $\mathcal{Y}$.
@@
$\begin{split}\ell_{\text{SegNCE}}(l,\mathcal{Z},i,\mathcal{Y})=\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \\ \frac{1}{\left\{j\middle|\begin{subarray}{c}j\in 1..|\mathcal{Z}|\\ y_{j}=y_{i}\\ i\neq j\end{subarray}\right\}}\sum_{\left\{j\middle|\begin{subarray}{c}j\in 1..|\mathcal{Z}|\\ y_{j}=y_{i}\\ i\neq j\end{subarray}\right\}}\ell_{\text{NCE}}(l,\mathbf{z}_{i},\mathcal{Z},j)\end{split}$ (6)
@@ III-D Model training
The total loss function for $G$ is given in equation 8. We include an identity PatchNCE loss following [29]. The full discriminator loss is given in Equation 9 and is simply the GAN objective.
@@ IV-A1 Training Details
Our generator $G$ is a 12M parameter ResNet-based network. The discriminator $D$ is a wider three-layer CNN with 11M parameters. Additionally, we parameterize the $H_{l}$ in Equations 3 and 5 as a 2-layer MLP with 700k parameters.
@@
We curate three test datasets: Fixed View, Randomized View, and Wrist View. Each test set contains 128 sim/real image pairs. The Fixed View test set contains sim and real images from the same fixed view as $\mathcal{D}_{B}$. The randomized view testset contains both sim and real viewpoints randomized within the box of size (100, 100, 84) cm. In order to create paired images for the Randomized View testset, we use the robot state in conjunction with AprilTags for coke can and camera pose estimation. Note that these are not needed for the deployment of MANGO, only for test set creation.
TABLE I: Sim2real Unpaired Image Translation FID Scores. Each score is the average score of two models trained with different random seeds. $D$ is the discriminator network and $\mathcal{D}_{A}$ is the sim training image dataset.
@@ IV-A2 Results
TABLE II: Sim2sim Experiment Results. Success rates are reported for 50 rollout of three random seeds. FID scores are calculated by comparing training data with the Oracle dataset.
@@
$\dagger$} Average pairwise LPIPS computed on 1000 randomly sampled images.
@@
We hypothesize that off-the-shelf methods like CUT and CycleGAN struggle on robot dataset due to their lack of diversity. Typically, these unpaired image-to-image translation methods are tested on computer vision benchmark datasets containing diverse images scraped from the internet. In comparison, robotics datasets contain limited diversity. To support this claim, we computed average pairwise LPIPS (a learned perceptual similarity metric) [48] on various image datasets and on our own dataset, shown in Table III. Our $\mathcal{D}_{B}$ shows the lowest score for diversity.
@@ IV-B Simulation Experiments
We evaluate MANGO on simulated tasks from Robomimic [27] and Mimicgen[26] and give the results in Table II. Specifically, we measure the FID scores of data generated with MANGO, as well as the success rates of behavioral-cloning policies trained with the data. Instead of sim2real, we create two visually disparate simulation environments and run sim2sim experiments. Example observations from each domain and MANGO translations are given in Figure 4. We benchmark MANGO against policies with single-camera observations, untranslated Domain A data, and VISTA [37]. To provide a fair comparison, we only train MANGO on image observations from the same camera viewpoints and tasks used by VISTA, and report metrics on the same six evaluation tasks. ”Unseen object” tasks contain objects not seen in training, ”shared object” contains objects seen during training but in different contexts, and ”cross-embodiment” tasks are seen performed by the Rethink Sawyer robot instead of Franka Panda. We train MANGO with fixed camera data from the six tasks used for testing in domain B, and varied camera data of the eight training tasks in domain A. This ensures our model does not see views from the test viewpoint distribution, aside from the single fixed view. MANGO-trained BC policies obtain the highest success rate on 5 out of 6 tasks.
@@ IV-C Real Robot Experiments
TABLE IV: Success rates for imitation learning policies across tasks, viewpoint shifts, and data augmentation methods. We note that the only method comparable to MANGO is VISTA, which leverages a 4.5B parameter pretrained model in contrast MANGO’s 35M parameters.
@@
We benchmark MANGO on four real-world robotic manipulation tasks: “pick up coke”, “stack cups”, “close laptop”, and “stack cups”. We train a single image translation model for all tasks, where $\mathcal{D}_{A}$ consists of 59,520 simulated observations and 35,294 fixed-camera real observations from tasks. Our real robot setup is shown in Figure 3.
@@ IV-C1 Policy Training and Evaluation Details
We train action chunking transformer (ACT) [50] policies on synthetic data generated by MANGO. The generated data is translated from our digital twin demonstrations which leverage the task and motion planner from RLBench [19]. All policies are cotrained with 150 human-teleoperated demos with fixed-camera and wrist camera observations. ACT was chosen to isolate the effects of our image data as it does not incorporate any pretraining or language conditioning. We train each ACT policy for 10k epochs with a chunk size of 20. Rollouts are done without temporal aggregation.
@@
We evaluate each trained policy on 10 variations per task on four different camera viewpoints. The first camera viewpoint is the fixed-camera which the real demonstrations and MANGO training data are collected from. The three shifted viewpoints, depicted in Figure 5, are aggregated into the “Shifted Cam” column in Table IV for brevity.
@@ IV-C2 Results
Results for MANGO-trained policies and baselines are given in Table IV. We observe that our method is necessary for viewpoint robustness, as compared to ACT models trained on fixed-cam human demonstrations only, which obtain the lowest success rates on shifted cameras. We also observe that our method is necessary to bridge the sim2real gap, since the policies trained on sim demos without translation perform consistently worse than MANGO.
@@ V Discussion and Conclusion
We train a novel image generation method on simulated and real robot data. We observe that with only fixed-camera real data, our novel SegNCE loss, discriminator design, and modified PatchNCE loss enable translation of simulated demos to highly-realistic novel views. MANGO-generated data improves the robustness of downstream imitation learning policies to camera shift, as demonstrated by greatly increased success rates six simulated and four real-world manipulation tasks. We observe that our method is superior for sim2real translation in this setting, beating all other image translation methods we tested.

### v2

@@ Abstract
Vision-based policies for robot manipulation have achieved significant recent success, but are still brittle to distribution shifts such as camera viewpoint variations. Robot demonstration data is scarce and often lacks appropriate variation in camera viewpoints. Simulation offers a way to collect robot demonstrations at scale with comprehensive coverage of different viewpoints, but presents a visual sim2real challenge. To bridge this gap, we propose MANGO – an unpaired image translation method with a novel segmentation-conditioned InfoNCE loss, a highly-regularized discriminator design, and a modified PatchNCE loss. We find that these elements are crucial for maintaining viewpoint consistency during sim2real translation. When training MANGO, we only require a small amount of fixed-camera data from the real world, but show that our method can generate diverse unseen viewpoints by translating simulated observations. In this domain, MANGO outperforms all other image translation methods we tested. Imitation-learning policies trained on data augmented by MANGO are able to achieve success rates as high as 60% on views that the non-augmented policy fails completely on. For more results, visit: https://www.jeremiahcoholich.com/mango.
@@ I Introduction
Figure 1: (a) Standard image translation methods fail to generalize to new viewpoints when trained on a fixed-viewpoint target domain dataset. (b) Our method, MANGO, enables realistic generation of unseen viewpoints, which are used to improve the robustness of downstream robot polices.
For example, many tabletop manipulation datasets employ fixed, third person cameras for observations [12, 39, 23, 41, 21, 6, 46], making downstream policies brittle to camera viewpoint shifts. Indeed, we have observed that when robot polices are trained on fixed-camera datasets, changes in camera viewpoint during deployment cause success rates to crater (Table IV). Cameras are often fixed during demo collection to ensure consistent visual observations, avoid repeated calibration with depth or motion capture sensors, or simply for convenience. Generalizing to truly unseen viewpoints is difficult because a change in viewpoint affects the entire scene, and the model must implicitly estimate the robot’s position relative to the new camera position.
@@ III-A Image-to-image Translation
We propose MANGO, a novel unpaired image-to-image translation method to translate visual observations from sim to real (Figure 2). MANGO is trained on a small, fixed-viewpoint dataset of real images yet is capable of accurately translating viewpoint-diverse observations from a simple digital twin to realistic unseen viewpoints.
The objective of unpaired image-to-image translation is to translate images from domain $A$ to domain $B$ without access to a paired dataset of images $\mathcal{D}_{paired}=\{d_{A},\;d_{B}|d_{A}\in A,\;d_{b}\in B\}_{i=0}^{N}$. Instead, we learn from two disjoint datasets $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$, where our domain $A$ is simulation, our domain $B$ is the real-world, and $|\mathcal{D}_{A}|>|\mathcal{D}_{B}|$. This problem is considered unsupervised because there is no label, or ground-truth image, in $\mathcal{D}_{B}$ that images in $\mathcal{D}_{A}$ map to.
@@ III-B Style Loss
One assumption underlying image-translation GANs, such as CycleGAN[51] and CUT [29], is that the shared attributes among all images in $\mathcal{D}_{B}$ constitute the target domain “style”. However, our real-world robot image observations in $\mathcal{D}_{B}$ only differ from one another in robot and object poses. Much of the image content, such as the background and tabletop, is nearly identical in all images in $\mathcal{D}_{B}$. A naïve discriminator will memorize the repetitive details and force the generator to recreate them. To mitigate this problem, Pix2pix [17] proposed a “PatchGAN”, where the discriminator only receives local image patches and cannot therefore enforce global image elements. We take this a step further and randomly sample patch locations and apply per-patch random rotations. This process is shown in Figure 2. The result is a highly-regularized discriminator capable of enforcing the style of $\mathcal{D}_{B}$ on images with viewpoints not seen in $\mathcal{D}_{B}$.
@@ III-C1 Modified PatchNCE Loss
The PatchNCE loss consists of an InfoNCE loss across encoder features generated by a source domain image $d_{A}\in D_{A}$ and its corresponding translated output image $G(d_{A})$. For an input image $d$, we randomly sample $N$ latent features from the encoder at $L$ different layers. We call the set of features at layer $l$ $\mathcal{Z}_{l}$ and $|\mathcal{Z}_{l}|=N\ \forall\ l\in\{l_{0},...,l_{L}\}$. The translated image $\hat{d}$ is passed through the encoder again to obtain $\hat{\mathcal{Z}}_{l}\ \forall\ l\in\{l_{0},...,l_{L}\}$. All $\hat{\mathcal{Z}}_{l}$ are obtained from the same feature map indices as in $\mathcal{Z}_{l}$.
The InfoNCE loss for feature $i$ in encoder layer $l$ is given by Equation 2. This is the categorical cross-entropy loss on the probability that a feature $\mathbf{z}\in\mathcal{Z}$ will be correctly classified as the corresponding feature in $\hat{\mathcal{Z}}$, based on a scoring function $\rho_{l}(\cdot)$. $\tau$ is a temperature hyperparameter.
@@
The reasoning behind Equation 2 is that input and output features from the same feature map locations are “positive” samples and should have high similarity scores. All other features are “negative” samples and should be repelled. However, we observe that many different input image patches are highly similar due to repeated patterns or textures in robot image datasets, which include background elements, the tabletop, etc. Furthermore, in the simulated image dataset $\mathcal{D}_{A}$, these regions all have identical pixel values due to simplistic rendering. Therefore, Equation 2 will repel many false negative features.
@@ III-C2 Segmentation NCE Loss
We leverage the simulator used to generate $\mathcal{D}_{A}$ to obtain ground-truth segmentation maps for each image. We propose an InfoNCE loss which clusters generator features by segmentation category in order to ensure that object boundaries are preserved during translation. We note that for object-centric manipulation, this is a crucial aspect to preserve.
Each image in $\mathcal{D}_{A}$ contains $C$ segmentation classes and each feature $\mathbf{z}_{i}\in\mathcal{Z}$ generated from $d\sim D_{A}$ has an associated class label $y_{i}\in\mathcal{Y}$. In the case when a layer’s feature map is of a lower resolution than the input image, we scale the image segmentation with nearest-neighbors downsampling to obtain $y_{i}$.
@@
$\begin{split}\ell_{\text{SegNCE}}(l,\mathcal{Z},i,\mathcal{Y})=\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \\ \frac{1}{\left|\left\{j\middle|\begin{subarray}{c}j\in 1..|\mathcal{Z}|\\ y_{j}=y_{i}\\ i\neq j\end{subarray}\right\}\right|}\sum_{\left\{j\middle|\begin{subarray}{c}j\in 1..|\mathcal{Z}|\\ y_{j}=y_{i}\\ i\neq j\end{subarray}\right\}}\ell_{\text{NCE}}(l,\mathbf{z}_{i},\mathcal{Z},j)\end{split}$ (6)
@@ III-D Model training
The total loss function for $G$ is given in equation 8. We include an identity PatchNCE loss for regularization following [29]. The full discriminator loss is given in Equation 9 and is simply the GAN objective.
@@ IV-A1 Training Details
Our generator $G$ is a 12M parameter ResNet-based network. The discriminator $D$ is a wider three-layer CNN with 11M parameters. Additionally, we parameterize the $H_{l}$ in Equations 3 and 5 as a two-layer MLP with 700k parameters.
@@
We curate three test datasets: Fixed View, Randomized View, and Wrist View. Each test set contains 128 sim/real image pairs. The Fixed View test set contains sim and real images from the same fixed view as $\mathcal{D}_{B}$. The randomized view testset contains both sim and real viewpoints taken from a camera placed randomly within a box of dimensions (100, 100, 84) cm. In order to create paired images for the Randomized View testset, we use the robot state in conjunction with AprilTags for coke can and camera pose estimation. Note that these are not needed for the deployment of MANGO, only for test set creation.
TABLE I: Sim2real Unpaired Image Translation FID Scores. Each score is averaged from two models trained with different seeds. $D$ is the discriminator and $\mathcal{D}_{A}$ is the sim training image dataset.
@@ IV-A2 Results
TABLE II: Sim2sim Experiment Results. Success rates are the average of 50 rollouts. FID scores are computed across target-domain generated images and target-domain oracle images.
@@
$\dagger$} Results computed on $10^{3}$ randomly sampled images.
@@
We hypothesize that off-the-shelf methods like CUT and CycleGAN struggle with robotics datasets due to their lack of diversity. Typically, unpaired image-to-image translation methods are benchmarked on computer vision datasets containing diverse images scraped from the internet. In comparison, robotics datasets may be collected from a single setup. To support this claim, we compute the average pairwise LPIPS (a learned perceptual distance metric) [48] on various image datasets as a measure of diversity. As shown in Table III, our $\mathcal{D}_{B}$ shows the lowest score.
@@ IV-B Simulation Experiments
We evaluate MANGO on simulated tasks from Robomimic [27] and Mimicgen[26] and give the results in Table II. Specifically, we measure the FID scores of data generated with MANGO, as well as the success rates of behavioral-cloning policies trained with the data. Instead of sim2real, we create two visually disparate simulation environments and run sim2sim experiments. Example observations from each domain and MANGO translations are given in Figure 4. We benchmark MANGO against policies trained on single-camera observations, untranslated Domain A data, and VISTA [37]. To provide a fair comparison, we only train MANGO on image observations from the same camera viewpoints and tasks used by VISTA, and report metrics on the same six evaluation tasks. “Unseen object” tasks contain objects not seen in training, “shared object” contains objects seen during training but in different contexts, and “cross-embodiment” tasks are seen performed by the Rethink Sawyer robot instead of Franka Panda. We train MANGO with fixed camera data from the six tasks used for testing in domain B, and varied camera data of the eight training tasks in domain A. This ensures our model does not see views from the test viewpoint distribution, aside from the single fixed view. MANGO-trained BC policies obtain the highest success rate on 5 out of 6 tasks.
@@ IV-C Real Robot Experiments
TABLE IV: Success rates for imitation learning policies across tasks, viewpoint shifts, and data augmentation methods. The only method comparable to MANGO is VISTA, which uses a 4.5B parameter pretrained model in contrast MANGO’s 35M parameters.
@@
We benchmark MANGO on four real-world robotic manipulation tasks: “pick up coke”, “stack cups”, “close laptop”, and “stack blocks”. We train a single image translation model for all tasks, where $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ consist of 59,520 simulated observations and 35,294 fixed-camera real observations from tasks. Our real robot setup is shown in Figure 3.
@@ IV-C1 Policy Training and Evaluation Details
We train action chunking transformer (ACT) [50] policies on synthetic data generated by MANGO. The generated data is translated from our digital twin demonstrations which leverage the task and motion planner from RLBench [19]. All policies are cotrained with 150 human-teleoperated demos with fixed-camera and wrist camera observations. ACT was chosen to isolate the effects of our generated image data as it does not incorporate any pretraining or language conditioning. We train each ACT policy for 10k epochs with a chunk size of 20. Rollouts are done without temporal aggregation.
@@
We evaluate each trained policy on 10 variations per task on four different camera viewpoints. The first camera viewpoint is the fixed-camera which the real demonstrations and MANGO training data are collected from. The three shifted viewpoints, depicted in Figure 5, are aggregated into the “Shifted Cam” column in Table IV.
@@ IV-C2 Results
Results for MANGO-trained policies and baselines are given in Table IV. With MANGO, we observe large increases in for viewpoint robustness as compared to models trained on fixed-camera human demonstrations only. We also observe that our method is necessary to bridge the sim2real gap, since the policies trained on sim demos without translation perform consistently worse than MANGO.
@@ V Discussion and Conclusion
We propose a novel image generation method, MANGO, which is trained on simulated and real robot data. We observe that with only fixed-camera real data, our novel SegNCE loss, discriminator design, and modified PatchNCE loss enable translation of unseen-viewpoint simulated observations to realistic real-world observations. MANGO-generated data improves the robustness of downstream imitation learning policies to camera shift, as demonstrated by greatly increased success rates six simulated and four real-world manipulation tasks. We observe that our method is superior for sim2real translation in this setting, beating all other image translation methods we tested.